OpenStudy (anonymous):

What are the solutions of the quadratic equation below? 2x^2 - 8x - 42 = 0

OpenStudy (anonymous):

have answer choices

OpenStudy (nurali):

2x^(2)-8x-42=0 Factor out the GCF of 2 from each term in the polynomial. 2(x^(2))+2(-4x)+2(-21)=0 2(x^(2)-4x-21)=0 2(x+3)(x-7)=0 x=-3, x=7

OpenStudy (anonymous):

yes it was, thankyaaa!

OpenStudy (nurali):

My Pleasure.